OpenStudy (anonymous):

Please help me it's urgent! The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HI!!!

OpenStudy (anonymous):

The perimeter of a rectangle is width + width + length + length So the perimeter is x + x + (x + 3) + (x + 3) Using perimeter = 58: x + x + (x + 3) + (x + 3) = 58 4x + 6 = 58 4x = 52 x = 13cm So the width = 13cm and the length = 16cm

OpenStudy (misty1212):

we need a variable what do you want to call the width?

OpenStudy (misty1212):

btw the answer above is very nice, but it is not the answer to your question

OpenStudy (misty1212):

name your variable and we can do it

OpenStudy (anonymous):

its 16

OpenStudy (anonymous):

x is fine :P And @Starbucks there isn't a 16 in my multiple choice.

OpenStudy (misty1212):

ok so we say the width is \(x\) and the length is 5 more than 3 times the width, so the length is \[3x+5\]

OpenStudy (misty1212):

then we add them up two of each we get \[x+x+3x+5+3x+5=58\] combining like terms gives \[8x+10=58\]

OpenStudy (anonymous):

my bad

OpenStudy (misty1212):

solve that one for \(x\) still one more step after that

OpenStudy (misty1212):

can you solve \[8x+10=58\]?

OpenStudy (anonymous):

6?

OpenStudy (misty1212):

yes but that is not the final answer

OpenStudy (anonymous):

Is it 23 inches? I did 3*6+5

OpenStudy (misty1212):

you are asked for the LENGTH and we solved for the width the length is \[3\times 6+5\]

OpenStudy (misty1212):

yeah 23

OpenStudy (anonymous):

Yayyyy!!! Omg thank you so much hahaha.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]